# Two-Sample Testing with Block-Wise Missingness in Multi-Source Data

### v1

@@ 2.1 Framework of Block-Pattern Enhanced Test (BPET)
Let $T=T(Z_{1}^{\mathrm{obs}},S_{1},\dots,Z_{N}^{\mathrm{obs}},S_{N};J_{1},\dots,J_{N})$ be a test statistic computed from the observed data $(Z_{i}^{\mathrm{obs}},S_{i})$ with group labels $J_{i}\in\{X,Y\}$.
(i) The standard permutation distribution of $T$ is obtained by recomputing $T$ after uniformly permuting the labels $\{J_{i}\}$ over all $\binom{N}{m}$ assignments with $m$ labels $X$ and $n$ labels $Y$, keeping $\{(Z_{i}^{\mathrm{obs}},S_{i})\}$ fixed. Under $H_{0}$, this is called the permutation null distribution.
@@
Let $\tilde{T}=T(Z_{1}^{\mathrm{obs}},S_{1},\dots,Z_{N}^{\mathrm{obs}},S_{N};\tilde{J}_{1},\dots,\tilde{J}_{N})$ with $\tilde{J}_{i}\in\{X,Y\}$ denote the statistic recomputed under permutation. The following theorem establishes the failure of the standard permutation scheme when the groups differ in their marginal pattern distributions, with proof given in Supplementary A.
Theorem 1 (Failure of standard permutation under unequal pattern distributions).
@@
but there exists at least one $s$ such that $P_{X}(S=s)\;\neq\;P_{Y}(S=s)$. Then under $H_{0}$, the distributions of $T$ and $\tilde{T}$ differ when $\{\tilde{J}_{i}\}$ are generated by the standard permutation scheme. Consequently, $p$-values based on the standard permutation distribution are invalid.
Theorem 1 demonstrates that standard permutation can fail whenever the two groups differ in their marginal distribution of missingness patterns, even if their conditional distributions within each pattern are identical. In such cases, $H_{0}$ may be rejected not because $F_{X}\neq F_{Y}$, but because of differences in the pattern distributions, leading to inflated type I error. This motivates restricting permutations to occur within patterns.
@@
Let $\tilde{T}$ denote the statistic computed using pattern-wise permutation. Under $H_{0}:F_{X}=F_{Y}$, a sufficient condition for $\tilde{T}\overset{d}{=}T$ is
$\frac{P_{X}(S=s\mid Z=z)}{P_{X}(S=s)}=\frac{P_{Y}(S=s\mid Z=z)}{P_{Y}(S=s)}\quad\text{ for all }s\text{ with }P_{g}(S=s)>0,\;g\in\{X,Y\}.$ (2)
This condition is strictly weaker than MCAR, which requires $P_{g}(S=s\mid Z=z)=P_{g}(S=s)\text{ a.s. }F_{g}\text{ for all }s$, $g\in\{X,Y\}$.
The proof is given in Supplementary B. Condition (2) is strictly weaker than both MCAR and the requirement for validity of standard permutation: whenever either of those holds, pattern-wise permutation is also valid. To our knowledge, this is the first general criterion that guarantees the validity of permutation inference under block-wise missingness. Under Condition (2), labels are exchangeable within each pattern, and the pattern-wise permutation distribution coincides with the null distribution conditional on the missingness patterns.
In the remainder of the paper we assume that Condition (2) holds. BPET can be combined with a wide range of two-sample tests; Sections 2.22.3 develop its implementation within a graph-based framework tailored to block-wise missing data.
2.2 NNG-induced Rank
In this section, we incorporate the RISE method with the proposed BPET framework. RISE method uses the rank of observations; in this paper, we focus on the $k$-nearest neighbor graph (NNG)-induced ranking scheme. We first introduce the $k$-NNG-based rank construction in RISE, then describe its adaptation to pattern-wise settings in BPET framework.
There exist a number of non-parametric methods for two-sample testing, each have their strength on different types of data. RISE stands out as a powerful graph-based two-sample testing approach, utilizing ranks induced from the $k$-NNG or $k$-minimum spanning tree (MST) to construct the test statistic, which is more flexible when dealing with multiple types of data (or multiple modalities) simultaneously (Zhou and Chen, 2023). For simplicity, we focus on $k$-NNG in this paper since the test can be easily adjusted to other similarity graphs with the same theoretical properties.
@@ 2.2 NNG-induced Rank
For any pair of observations $(Z_{i},Z_{j})$, compute their dissimilarity using (1) or other admissible metrics. The graph-induced rank of $(Z_{i},Z_{j})$ is defined as
@@
if $Z_{j}$ is the $k^{\prime}$th nearest neighbor of $Z_{i}$ for $1\leq k^{\prime}\leq k$ and $\mathrm{rank}(Z_{j}\mid Z_{i})=0$ otherwise.
When ties occur in the dissimilarity values, the tied neighbors are assigned the average of the ranks they would otherwise occupy.
Now, we consider how to adapt RISE into the proposed BPET framework; the key idea is to define a systematic ranking scheme for any pattern-pairs. First, we define the (pattern) pair-wise $k$-NNG. Given two patterns $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$, the pair-wise $k$-NNG is constructed based on observations from $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, where each observation in $\mathcal{Z}^{(\alpha)}$ is connected to its $k$ nearest neighbors in $\mathcal{Z}^{(\beta)}$. An illustration of such pair-wise NNGs is shown in Figure 3. It is noteworthy that the structure of the $k$-NNG depends on the relationship between $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$:
  • If $\alpha=\beta$, then the $k$-NNG is a standard $k$-NNG, where edges may connect any pair of observations within the same pattern as depicted in the left panel of Figure 3.
  • If $\alpha\neq\beta$, then the $k$-NNG is a bipartite graph, each observation in $\mathcal{Z}^{(\alpha)}$ connects to its nearest neighbors in $\mathcal{Z}^{(\beta)}$, with no within-pattern edges as depicted in the right panel.
@@
With the ranking scheme in Definition 2, the rank of an observation $Z_{j}\in\mathcal{Z}^{(\beta)}$ with respect to $Z_{i}\in\mathcal{Z}^{(\alpha)}$ is denoted by $R_{ij}$. The total rank matrix is then given by $\mathbf{R}=[R_{ij}]_{i,j=1}^{N}$. Grouping observations by missingness pattern, we denote $\mathbf{R}(\mathcal{W}^{(\alpha)},\mathcal{W}^{(\beta)})=[R_{ij}:Z_{i}\in\mathcal{W}^{(\alpha)},\;Z_{j}\in\mathcal{W}^{(\beta)}],\,\mathcal{W}\in\{\mathcal{X},\mathcal{Y},\mathcal{Z}\}.$ Then for $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, the corresponding pair-wise rank matrix is defined as
$\mathbf{R}^{(\alpha\beta)}=\mathbf{R}(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})=\begin{bmatrix}\mathbf{R}(\mathcal{X}^{(\alpha)},\mathcal{X}^{(\beta)})&\mathbf{R}(\mathcal{X}^{(\alpha)},\mathcal{Y}^{(\beta)})\\ \mathbf{R}(\mathcal{Y}^{(\alpha)},\mathcal{X}^{(\beta)})&\mathbf{R}(\mathcal{Y}^{(\alpha)},\mathcal{Y}^{(\beta)})\end{bmatrix}=[R_{ij}^{(\alpha\beta)}]_{i,j=1}^{N_{\alpha},N_{\beta}}.$
@@ 2.3 Test Statistics
We now describe how to aggregate the pattern-pair level rank information into global test statistics. For simplicity, we adopt a symmetrized version of the rank matrix by setting $\mathbf{R}\leftarrow\tfrac{1}{2}(\mathbf{R}+\mathbf{R}^{\top})$, which guarantees $R_{ij}^{(\alpha\beta)}=R_{ji}^{(\beta\alpha)}$. Throughout, $\mathbf{R}$ denotes this symmetrized rank matrix.
For each pattern pair $(\alpha,\beta)$, define
@@
which represent the within-sample rank sums for groups $X$ and $Y$. When $\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}=0$, dissimilarities are set to zero and the resulting ranks carry no information; such pairs are therefore excluded from the construction of test statistics. By symmetry, $U_{x}^{(\alpha\beta)}=U_{x}^{(\beta\alpha)}$ and $U_{y}^{(\alpha\beta)}=U_{y}^{(\beta\alpha)}$.
Before defining the test statistics, we examine the behavior of $U_{x}^{(\alpha\beta)}$ and $U_{y}^{(\alpha\beta)}$ in an illustrative example. Consider $m=n=50$ with two sources, each $50$-dimensional, generated independently from Gaussian distributions. Block-wise missingness is introduced by removing each source independently with probability $0.5$ (regenerating if both are missing), yielding three observed patterns. The $X$-samples are drawn from $\mathcal{N}_{100}(\mathbf{0}_{100},\mathbf{I}_{100})$, while the $Y$-samples follow one of four distributions: (a) the same null distribution; (b) a location shift $\mathcal{N}_{100}(\mathbf{1}_{100},\mathbf{I}_{100})$; (c) a scale shift $\mathcal{N}_{100}(\mathbf{0}_{100},4\mathbf{I}_{100})$; or (d) a mixed alternative $\mathcal{N}_{100}(0.5\mathbf{1}_{100},2\mathbf{I}_{100})$.
Figure 4: Heatmaps of pattern-wise rank matrices from $10$-NNGs (three patterns).
Figure 4 shows heatmaps of the rank matrices. Under the location alternative, both $U_{x}^{(\alpha\beta)}$ and $U_{y}^{(\alpha\beta)}$ increase relative to their null values, whereas under the scale alternative, one tends to increase while the other decreases. For the mixed alternative, both $U_{x}^{(\alpha\beta)}$ and $U_{y}^{(\alpha\beta)}$ deviate from the null in different directions. These patterns motivate two complementary aggregation strategies: a vectorized form that is sensitive to localized deviations across patterns, and a low-dimensional aggregated form that captures overall distributional shifts.
Let $\mathcal{I}=\{(\alpha,\beta):1\leq\alpha\leq\beta\leq n_{P},\;\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}>0\}$ be the set of valid pattern pairs. Define the vector $V$ as the concatenation of
$\big{(}U_{x}^{(\alpha\beta)}-\mu_{x}^{(\alpha\beta)},\;U_{y}^{(\alpha\beta)}-\mu_{y}^{(\alpha\beta)}\big{)},\quad(\alpha,\beta)\in\mathcal{I},$
so that $V\in\mathbb{R}^{2|\mathcal{I}|}$. Here $\mu_{x}^{(\alpha\beta)}=\mathbb{E}(U_{x}^{(\alpha\beta)})$ and $\mu_{y}^{(\alpha\beta)}=\mathbb{E}(U_{y}^{(\alpha\beta)})$, with expectations taken under the pattern-wise permutation null distribution. The vectorized Block-wise RISE statistic (BRISE-v) is
@@
where $\boldsymbol{\Sigma}_{v}=\mathrm{Cov}(V)$ is also evaluated under the pattern-wise permutation null distribution. By construction, $T_{v}$ increases when any $U_{x}^{(\alpha\beta)}$ or $U_{y}^{(\alpha\beta)}$ deviates from its null expectation, making it sensitive to a broad range of alternatives.
Alternatively, consider the aggregated sums
${U}_{x}=\sum_{(\alpha,\beta)\in\mathcal{I}}U_{x}^{(\alpha\beta)}=\sum_{i=1}^{m}\sum_{j=1}^{m}R_{ij}\quad\mathrm{and}\quad{U}_{y}=\sum_{(\alpha,\beta)\in\mathcal{I}}U_{y}^{(\alpha\beta)}=\sum_{i=m+1}^{N}\sum_{j=m+1}^{N}R_{ij},$
which are equivalent to summing over all within-group ranks. The congregated Block-wise RISE statistic (BRISE-c) is
$T_{c}={\bar{U}}^{\top}\boldsymbol{\Sigma}_{c}^{-1}{\bar{U}},$ (4)
where $\bar{U}=({U}_{x}-\mu_{x},{U}_{y}-\mu_{y})^{\top}$, $\mu_{x}=\mathbb{E}({U}_{x})$, $\mu_{y}=\mathbb{E}({U}_{y})$, and $\boldsymbol{\Sigma}_{c}=\mathrm{Cov}(({U}_{x},{U}_{y})^{\top})$.
BRISE-v is sensitive to heterogeneous local deviations but requires stable estimation of $\boldsymbol{\Sigma}_{v}$, while BRISE-c is computationally more stable and better suited when $n_{P}$ is large or sample sizes are limited.
@@
$\tilde{U}_{x}=\sum_{(\alpha,\beta)\in\mathcal{I}}a_{x}^{(\alpha\beta)}U_{x}^{(\alpha\beta)},\qquad\tilde{U}_{y}=\sum_{(\alpha,\beta)\in\mathcal{I}}a_{y}^{(\alpha\beta)}U_{y}^{(\alpha\beta)}$
@@
To avoid instability from extremely rare patterns, it is advisable to discard $\mathcal{P}_{\alpha}$ with very small sizes, for instance those with $\min\{m_{\alpha},n_{\alpha}\}<n_{\mathrm{thres}}$ or $N_{\alpha}<p_{\mathrm{thres}}\max_{\beta}N_{\beta}$, using defaults such as $n_{\mathrm{thres}}=2$ and $p_{\mathrm{thres}}=0.1$. The covariance matrices $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$ should also be invertible. Simulations in Supplementary M show that both are typically stable, with $\boldsymbol{\Sigma}_{c}$ more robust in small samples. In rare cases of ill-conditioning, which occurs when some quantities $U_{g}^{(\alpha\beta)}$, $g\in\{x,y\}$, are linearly dependent, a subset of these terms can be removed to ensure invertibility of the covariance matrix.

### v2

@@ 2.2 A Pattern-wise Permutation Procedure
Let $T=T(Z_{1}^{\mathrm{obs}},S_{1},\dots,Z_{N}^{\mathrm{obs}},S_{N};G_{1},\dots,G_{N})$ be a test statistic computed from the observed data $(Z_{i}^{\mathrm{obs}},S_{i})$ with group labels $G_{i}\in\{X,Y\}$.
(i) The standard permutation distribution of $T$ is obtained by recomputing $T$ after uniformly permuting the labels $\{G_{i}\}$ over all $\binom{N}{m}$ assignments with $m$ labels $X$ and $n$ labels $Y$, keeping $\{(Z_{i}^{\mathrm{obs}},S_{i})\}$ fixed. Under $H_{0}$, this is called the permutation null distribution.
@@
As an illustration of Definition 1(ii), suppose there are two missingness patterns $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. If $\mathcal{P}_{1}$ contains $5$ observations with $3$ from Group $X$ and $2$ from Group $Y$, then the labels are permuted uniformly over all assignments preserving these counts. The same procedure is applied independently within $\mathcal{P}_{2}$, preserving $(m_{\alpha},n_{\alpha})$ for each pattern.
The proposed pattern-wise permutation is closely related to classical stratified permutation tests, where labels are permuted within predefined strata (Fisher, 1935; Lehmann and Romano, 2005). In our setting, the strata are defined by missingness patterns rather than fully observed covariates. Unlike classical stratification, missingness patterns may depend on unobserved components of the data, and the validity of such permutation schemes under block-wise missingness has not been previously established. In this work, we provide a sufficient condition (Theorem 2) that guarantees the validity of pattern-wise permutation and, consequently, the proposed framework.
2.3 Validity under Different Missing Assumptions
We examine the validity of standard permutation under block-wise missingness and provide a sufficient condition under which pattern-wise permutation yields valid inference. We further show that this condition is strictly weaker than MCAR and can accommodate MNAR mechanisms.
Let $\widetilde{T}=T(Z_{1}^{\mathrm{obs}},S_{1},\dots,Z_{N}^{\mathrm{obs}},S_{N};\widetilde{G}_{1},\dots,\widetilde{G}_{N})$ with $\widetilde{G}_{i}\in\{X,Y\}$ denote the statistic recomputed under permutation. The following theorem establishes the failure of the standard permutation scheme when two groups/samples differ in their marginal missing pattern distributions.
Theorem 1 (Failure of standard permutation under unequal missing pattern distributions).
@@ 2.3 Validity under Different Missing Assumptions
but there exists at least one $s$ such that $P_{S\mid G=X}(s)\;\neq\;P_{S\mid G=Y}(s)$. Then under $H_{0}$, the distributions of $T$ and $\widetilde{T}$ differ when $\{\widetilde{G}_{i}\}$ are generated by the standard permutation scheme. Consequently, $p$-values based on the standard permutation distribution are invalid.
Theorem 1 shows that standard permutation can fail whenever the two groups differ in their marginal distributions of missingness patterns, even if their conditional distributions within each pattern are identical. In such cases, rejection of $H_{0}$ may be driven by differences in missingness patterns rather than distributional differences between $F_{X}$ and $F_{Y}$, leading to inflated type-I error. Since unequal pattern distributions commonly arise in block-wise missing data due to heterogeneous data collection processes, standard permutation may be inappropriate in this setting.
@@
Let $\widetilde{T}$ denote the statistic computed using pattern-wise permutation. Under $H_{0}:F_{X}=F_{Y}$, a sufficient condition for $\widetilde{T}\overset{d}{=}T$ is
$\frac{P_{S\mid Z,G=X}(s\mid z)}{P_{S|G=X}(s)}=\frac{P_{S\mid Z,G=Y}(s\mid z)}{P_{S|G=Y}(s)}\quad\text{ for all }z,s\text{ with }P_{S|G}(s):=P_{G}(S=s)>0,$ (2)
where $G\in\{X,Y\}$. This condition is strictly weaker than MCAR, which requires $P_{S\mid Z,G}(s\mid z)=P_{S\mid G}(s)\text{ a.s. }F_{G}\text{ for all }s$, $G\in\{X,Y\}$.
Condition (2) is strictly weaker than MCAR and the assumptions required for standard permutation; whenever either holds, pattern-wise permutation is valid. Importantly, Condition (2) allows the missingness mechanism to depend on the full data vector, including unobserved components, and thus accommodates MNAR mechanisms. This condition differs fundamentally from the MAR assumption, which restricts dependence to observed data only. Instead, Condition (2) requires that the dependence of the missingness mechanism on the data is proportional across groups. To our knowledge, this is the first general criterion guaranteeing valid permutation inference under block-wise missingness.
Under Condition (2), labels are exchangeable within each missingness pattern, and the pattern-wise permutation distribution coincides with the null distribution conditional on the observed patterns. Throughout the remainder of the paper, we assume that Condition (2) holds.
3 A Graph-Induced Rank-Based Test within BPET
In this section, we incorporate the rank-weighted similarity graph method RISE into the proposed BPET framework to integrate modalities of heterogeneous types. RISE is a flexible graph-based two-sample testing approach that constructs test statistics using ranks induced from similarity graphs, such as the $k$-nearest neighbor graph ($k$-NNG) or the $k$-minimum spanning tree ($k$-MST), and naturally accommodates both Euclidean and non-Euclidean data geometries (Zhou and Chen, 2023). For simplicity, we focus on the $k$-NNG in this paper and extend the $k$-NNG-induced ranking scheme to block-wise missing data via pattern-aware graphs.
3.1 Graph-induced Rank
For a given $k$, the $k$-NNG is a directed graph in which each node represents an observation and each directed edge points from an observation to its $k$ nearest neighbors under a chosen dissimilarity measure. Based on the $k$-NNG, we define the induced rank as follows.
@@ 3.1 Graph-induced Rank
For any pair of observations $(Z_{i},Z_{j})$, the graph-induced rank of $(Z_{i},Z_{j})$ is defined as
@@
if $Z_{j}$ is the $k^{\prime}$th nearest neighbor of $Z_{i}$ for $1\leq k^{\prime}\leq k$ and $\mathrm{rank}(Z_{j}\mid Z_{i})=0$ otherwise. When ties occur in the dissimilarity values, the tied neighbors are assigned the average of the ranks they would otherwise occupy.
Now, we consider how to extend the $k$-NNG-induced rank to block-wise missing data. The key idea is to define a systematic ranking scheme that depends on the missingness patterns of each pair of observations. We first define pattern-aware $k$-NNGs. Given two observations $(Z_{i},Z_{j})$ belonging to missingness patterns $\mathcal{P}_{\alpha}$ and $\mathcal{P}_{\beta}$, respectively, the pattern-aware $k$-NNG is constructed using only observations from $\mathcal{Z}^{(\alpha)}\cup\mathcal{Z}^{(\beta)}$. In this graph, each node represents an observation in $\mathcal{Z}^{(\alpha)}$ or $\mathcal{Z}^{(\beta)}$, and each observation in $\mathcal{Z}^{(\alpha)}$ points to its $k$ nearest neighbors in $\mathcal{Z}^{(\beta)}$ via directed edges. An illustration is shown in Figure 3. The structure of the pattern-aware $k$-NNG depends on whether $\mathcal{P}_{\alpha}=\mathcal{P}_{\beta}$:
  • If $\alpha=\beta$, then the pattern-aware $k$-NNG is a standard $k$-NNG with only observation from $\mathcal{P}_{\alpha}$ or $\mathcal{P}_{\beta}$, where bi-directional edges are possible (see the left panel in Figure 3);
  • If $\alpha\neq\beta$, then the pattern-aware $k$-NNG is a bipartite graph, each observation in $\mathcal{Z}^{(\alpha)}$ points to its nearest neighbors in $\mathcal{Z}^{(\beta)}$, with no within-pattern edges (see the right panel in Figure 3).
Then, for any pair of $(Z_{i},Z_{j})$ from missing patterns $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, the rank of an observation $Z_{j}$ w.r.t. $Z_{i}$ is the $k$-NNG-induced rank based on the pattern-aware $k$-NNG.
@@
With the ranking scheme above, the rank of an observation $Z_{j}\in\mathcal{Z}^{(\beta)}$ with respect to $Z_{i}\in\mathcal{Z}^{(\alpha)}$ is denoted by $R_{ij}$. The corresponding total rank matrix is $\mathbf{R}=[R_{ij}]_{i,j=1}^{N}$. Denote $\mathbf{R}(\mathcal{W}^{(\alpha)},\mathcal{W}^{(\beta)})=[R_{ij}:Z_{i}\in\mathcal{W}^{(\alpha)},\;Z_{j}\in\mathcal{W}^{(\beta)}],\,\mathcal{W}\in\{\mathcal{X},\mathcal{Y},\mathcal{Z}\}.$ We can rearrange $\mathbf{R}$ into a block-wise matrix with each block corresponding to a pair of missing patterns within two samples/groups; for $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, the corresponding pattern-pair rank matrix (a sub-matrix of $\mathbf{R}$) is
$\mathbf{R}^{(\alpha\beta)}=\mathbf{R}(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})=[R_{ij}^{(\alpha\beta)}]_{i,j=1}^{N_{\alpha},N_{\beta}}:=\begin{bmatrix}\mathbf{R}(\mathcal{X}^{(\alpha)},\mathcal{X}^{(\beta)})&\mathbf{R}(\mathcal{X}^{(\alpha)},\mathcal{Y}^{(\beta)})\\ \mathbf{R}(\mathcal{Y}^{(\alpha)},\mathcal{X}^{(\beta)})&\mathbf{R}(\mathcal{Y}^{(\alpha)},\mathcal{Y}^{(\beta)})\end{bmatrix}$.
@@ 3.2 Test Statistics
We describe how to aggregate the pattern-pair rank matrices into global test statistics. For simplicity, we work with a symmetrized rank matrix $\mathbf{R}\leftarrow\tfrac{1}{2}(\mathbf{R}+\mathbf{R}^{\top})$, which guarantees $R_{ij}^{(\alpha\beta)}=R_{ji}^{(\beta\alpha)}$. Throughout, $\mathbf{R}$ denotes this symmetrized matrix. For each pair of missing patterns $(\mathcal{Z}^{(\alpha)},\mathcal{Z}^{(\beta)})$, define
@@
which represent the within-sample rank sums for Groups $X$ and $Y$.
An illustrative example in Supplementary L shows that different alternatives may induce different deviation patterns: under scale or mixed alternatives, $U_{x}^{(\alpha\beta)}$ and $U_{y}^{(\alpha\beta)}$ can deviate in opposite directions for specific pattern pairs, whereas under location alternatives, they tend to shift in the same direction, with deviations accumulating across patterns.
Motivated by this observation, we propose two complementary aggregation strategies: a vectorized statistic sensitive to localized pattern-specific deviations, and a low-dimensional aggregated statistic capturing overall distributional shifts. Let
$\mathcal{I}=\{(\alpha,\beta):1\leq\alpha\leq\beta\leq n_{P},\;\mathcal{P}_{\alpha}\mathcal{P}_{\beta}^{\top}>0\}.$
Define $V=\big(U_{x}^{(\alpha\beta)}-\mu_{x}^{(\alpha\beta)},\;U_{y}^{(\alpha\beta)}-\mu_{y}^{(\alpha\beta)}\big)^{\top}_{(\alpha,\beta)\in\mathcal{I}}\in\mathbb{R}^{2|\mathcal{I}|},$ where $\mu_{x}^{(\alpha\beta)}=\mathbb{E}(U_{x}^{(\alpha\beta)})$ and $\mu_{y}^{(\alpha\beta)}=\mathbb{E}(U_{y}^{(\alpha\beta)})$ are taken under the pattern-wise permutation null distribution. The vectorized Block-wise RISE statistic (BRISE-v) is
@@
with $\boldsymbol{\Sigma}_{v}=\mathrm{Cov}(V)$ under the same null. By construction, $T_{v}$ is sensitive to deviations arising from specific missing-pattern pairs.
Alternatively, define the aggregated rank sums
${U}_{x}=\sum_{(\alpha,\beta)\in\mathcal{I}}U_{x}^{(\alpha\beta)}=\sum_{i=1}^{m}\sum_{j=1}^{m}R_{ij},\qquad{U}_{y}=\sum_{(\alpha,\beta)\in\mathcal{I}}U_{y}^{(\alpha\beta)}=\sum_{i=m+1}^{N}\sum_{j=m+1}^{N}R_{ij},$
which correspond to summing all within-group ranks. The congregated Block-wise RISE statistic (BRISE-c) is
$T_{c}={\widebar{U}}^{\top}\boldsymbol{\Sigma}_{c}^{-1}{\widebar{U}},$ (4)
where $\widebar{U}=({U}_{x}-\mu_{x},{U}_{y}-\mu_{y})^{\top}$, $\mu_{x}=\mathbb{E}({U}_{x})$, $\mu_{y}=\mathbb{E}({U}_{y})$, and $\boldsymbol{\Sigma}_{c}=\mathrm{Cov}(({U}_{x},{U}_{y})^{\top})$. Compared to BRISE-v, BRISE-c aggregates information across all patterns and is therefore more sensitive to global distributional shifts.
The two statistics are complementary: BRISE-v targets heterogeneous local deviations across patterns, while BRISE-c emphasizes overall deviations. To avoid instability from extremely rare patterns when computing BRISE-v, we discard patterns with $\min\{m_{\alpha},n_{\alpha}\}<n_{\mathrm{thres}}$ or $N_{\alpha}<p_{\mathrm{thres}}\max_{\beta}N_{\beta}$. Throughout, we use $n_{\mathrm{thres}}=2$ and $p_{\mathrm{thres}}=0.1$.
@@
$\widetilde{U}_{x}=\sum_{(\alpha,\beta)\in\mathcal{I}}a_{x}^{(\alpha\beta)}U_{x}^{(\alpha\beta)},\qquad\widetilde{U}_{y}=\sum_{(\alpha,\beta)\in\mathcal{I}}a_{y}^{(\alpha\beta)}U_{y}^{(\alpha\beta)}$
@@
The covariance matrices $\boldsymbol{\Sigma}_{v}$ and $\boldsymbol{\Sigma}_{c}$ should be invertible to ensure $T_{v}$ and $T_{c}$ are well-defined. Simulations in Supplementary N show that both are typically stable, with $\boldsymbol{\Sigma}_{c}$ more robust in small samples. In rare cases of ill-conditioning, which occurs when some quantities $U_{g}^{(\alpha\beta)}$, $g\in\{x,y\}$, are linearly dependent, a subset of these terms can be removed to ensure invertibility of the covariance matrix.